FEL Upgrade Project Weekly Brief

November 11-15, 2002

We have experienced another extremely busy and productive week for the FEL Upgrade with installation and commissioning activities. Highlights include the following:

(1) all four of the large dipole magnets (GW’s) for the first optical chicane were qualified by the magnetic measurement crew, found to meet our demanding specifications (1/1000), and released for installation. Two have been installed and the second two will be installed next week; (2) we have identified several unique work-arounds for the late delivery of the recirculation hardware that allows our commissioning schedule to proceed (see WBS 9 write-up); with these work-arounds all beam transport for the first arc is now on site and being qualified for installation; (3) the optical cavity vacuum assemblies were installed in the final positions fore and aft of the wiggler; the High Voltage Power Supply tank for the electron gun was successfully pressure tested for operation.

Today an article was published in Nature by Gwyn Williams et al that describes our successful measurement and analysis of the high power (20 watt) , short pulse, terahertz radiation with IR Demo last year just before we shut down the device for FEL Upgrade installation.

The article is available at:
and an accompanying "News and Views" story is available at:

Effort was spent this week to analyze and document the FEL commissioning technical and cost status for presentation to Gil Graff at the ONR Program Office. Several scenarios were presented that prioritized FY03 FEL Upgrade activities with respect to potential FY03 funding levels. These scenarios will be reviewed at next week’s FEL Upgrade Semiannual Review for the review committee’s review and comment. We appreciate the time that was devoted to this task by Gil Graf at ONR and Wayne Skinner at our DOE Site Office.

The next Semiannual Project Review is next week on Wed. and Thurs., November 20-21. WBS managers spent some time this week in between their installation, check-out and commissioning activities making preparations for the review. We are all looking forward to showing off the new machine next week to our review committee.

George Neil and Steve Benson gave several presentations on the status of the FEL Upgrade program and scaling to higher power FEL systems at this week’s Directed Energy Professional Society Symposium hosted by NAVSEA and the Naval Postgraduate School at Monterey, CA.

FELs had very high visibility at the DEPS conference and in addition to George and Steve's papers, FEL Dept. members were listed as co-authors on three AES papers and two NPS papers.

WBS 3 (Beam Physics):
Results for various magnet measurements have been reviewed and the magnet’s performance certified.

Tune-up procedures are being simulated so as to provide final dump line layout specifications. An amusing result has emerged: Conservation of energy limits the amount of energy you can recover. (Corollary: If the bunch length is longer than 360 degrees, you can't recover it very well...!)

WBS 4 (Injector):
The door operating bellows flange was removed from the gun and successfully welded. The gun was pumped down, all ion pumps started and a satisfactory leak check was accomplished. A small amount of contamination was discovered in the gun and injector line and a hot N2 purge is being planned to remove the bulk of it prior to the bake. The extractor cart was removed from the ceramic stack and the corona rings and support rods were fit up. The SF6 tank is being readied for fit up with the high voltage tank so we can mark the operating probe position and complete fabrication of the resistor.

Gun HVPS - The Conditioning Resistor assembly is together and will be positioned to mark the location for its probe connector. The connector will then be welded into the aluminum shell. The HVPS Tank was also leak tested and passed this week.

WBS 6 (RF):
Quarter HVPS - The wiring for Unit #4 HVPS is complete and will be tested for PSS Certification noon today. The klystron cart is wired, but still lacks some LCW fittings. These fittings are on order. The software appears to be working fine in all aspects. The RF systems will be connected to the Quarter cavities early next week for SRF testing.

Zone 2 - Complete and operational. SRF tested to its Drive High values without problems.

Zone 4 - Installation is complete. SRF has tested the zone, but not yet to its Drive High values. They will check the cavity Q's and download files as the klystrons forward power seem more than adequate to produce the Drive High gradients. It will be checked again early next week.

WBS 8 (Instrumentation):
Drive Laser Controls - The acromag card is fully implemented into iocfel10 and all functions (Digital-I/O, ADC & DAC channels) are fully checked-out. The difficulty in bringing this card online was due to an oversight on our part in properly configuring the VME backplane to handle the interrupts (a lesson we won't soon forget). As a result, we now know why we had so much trouble with the OMS stepper-motor controller. Having resolved all our hardware issues with these two plug-in modules, their respective control elements (water pump, remote key status, shutter control, EO Cell bias control, SHG readback and remote/variable attenuation are all online. We are eager to log some operation time on these items now.

Drive Laser Pulse Controller (DLPC) Upgrade - The SSMC-RF connectors for the CO305 VME Module came in sooner than expected and so the board is being finished for testing next week. Also, as part of the DLPC upgrade, a full system print of the drive laser as a sub-system is being produced. This will document everything in the drive laser clean room from an as-built/controls perspective. With the basic connectivity of the drive laser's RF controls documented including the custom modifications to the Modelocker chassis, we are in a good position to lock-in the design of the automatic switch to internal RF as a part of the external RF interlock upgrade. This control function prevents the possibility of a "Q-Switch" and cooking the Second Harmonic Generation (SHG) crystal. The modified Modelocker chassis with the manual RF switch is this weeks picture link at: http://laser.jlab.org/wbs8/

Beam Loss Monitors (BLMs) - The F0151 module has been pulled from iocfel8 and given to EES as a the functional prototype to aid in the assembly of the rest of the boards. We will get it back as they finish new boards one at a time versus assembling the lot in parallel.

The new beam viewer design is complete and being manufactured (THANKS Casy!). This design is a 1"x4" silicon paddle fixed to a 6" conflat blank. The paddle has registration marks at 5mm with a major mark on center. There are two of these being built one for the strait ahead spectrometer and the second for the GW beam diverter downstream of the wiggler. All parts are on order for the Synchrotron Light Monitor (SLM) attenuators. This is a very simple design based on 2"x4" box aluminum beam. Assembly of the prototype should take place next week.

The old OPCHIC DC power supply has been reinstalled for powering the DW & DY (work around). It was moved into place and had its AC power connected. The FEL I&C team has wired into the PSS (they are much to busy) to enable certification next week during the scheduled PSS certification of the 1/4CM HPA. This is the LAST PSS connected power supply. The software group has begun work on the software and screens to support the old OPCHIC power supply. DC power to the 0F region corrector magnets is near completion. The new GW does not require a PSS connection since it has the leads covered and is only a 5 kW supply.

Those (2ga) leads will be run in next week. The optical cavity wiring (controls) is progressing, much of this activity is done after 1700 and must be coordinated with the (nearly) nightly lockups for RF commissioning. The control screen for the 4 channel BPMs are being worked, much of the hardware will be checked out next week.

Beamviewers for the wiggler have been installed and checkout is in progress. The 25 pin MPS breakout modules have been received from the vendor and are being populated in house. Beamviewer MPS cables were pulled into FL02B10 and termination is in progress. Work progressed on the BLM high voltage cable terminations.

Schematic capture for the HV Interlock digital I/O board have been received from EECAD and under review. Work has progressed on the system print for the HV interlock. The software for the Machine Protection System (MPS) is progressing very well. There are some screens done with the new cards all reading back. Many of the input cables are complete and ready to connect, this will be done as priorities allow.

WBS 9 (Transport):
Work Arounds - Work arounds are now implemented to compensate for the recirculation magnets that are late while maintaining the commissioning schedule. Dave Douglas has blessed all these efforts as worthy of pursuit.

1. To commission the linac without the first arc complete in December, we need a dipole to send the beam off at an angle at the end of the linac to phase the cavities. We installed an IR Demo Optical Chicane Dipole and its chamber just after the GX position with an innovative silicon wafer beam viewer developed by Kevin Jordan. It will be powered by the spare 66 kW supply in series with the DY below. Control was simplified by powering both provisional magnets on the same supply because machine commissioning uses only one of these magnets at any one time.

2. For first light commissioning using pulsed beam, mid January, use an IR Demo 180° Dipole (DY) to turn the first arc (along with the GQ and the GX dipoles that will be measured by then). Fortunately DY is rated for 80 MeV/c (the commissioning momentum) and its 2 inch aperture is adequate for the requirements of the first arc. The DY only needs new flanges on its chamber to adapt to the GQ-GX chambers.

We will extract pulsed electron beam using a GW dipole, three Quadrupoles past the optical klystron so we don’t hit the mirror with the electron beam. At that position, we can use an IR Demo optical chicane chamber for the 20° bend-out and we can easily stack SEG (steel shielding) blocks to absorb the beam. The GW dipole will be one of the dipoles intended for the UV Line. We will power it with the already installed power supply for the extraction chicane, (which is not powered during this phase of the effort). 3. For recirculation and high power lasing, starting near the end of February, install the 1st 180° dipole and GQs & GXs in the second arc with the IR Demo’s DY remaining in the first Arc and remove the GW above.

4. Since the Sextupoles are designed but not procured, work around using IR Demo SC Sextupoles. We are expanding them with shim blocks and pole shifts to fit over larger chamber and they are just strong enough for the commissioning energy of 80 MeV if we double their current and cool

Optical Chicane Dipoles (GW)
• Continued magnetic success. Magnetic tests on the stepper stand qualified all four GWs for the IR Machine. They meet David Douglas’ specifications for core field flatness in the transverse (± 1 part in 10,000), field integral over the good field region (± 1 part in 10,000) and exceeded the spec by a factor of two for core field and field integral within the magnet family (within a part in 1000) at the nominal machine energy of 145 MeV. Two were installed this week with the second two to follow next week.
• Additional tests verified the levels of flatness on the "Dipole" Stand where a loop-coil tests field integral flatness. The Group is using these magnets to calibrate the stand for use on the GQ-GX styles, coming next.
• Kudos to Magnet Enterprises International and the collection of machining and grinding vendors in the Oakland industrial area for making these magnets so beautifully, especially gluing the Purcell gap material down to the pole tips to the required precision.
• Kudos also to the Magnetic Measurement Group and their help from the Jlab Machine Shop for running a 16 hr. schedule and measuring the 4 GWs all in one week, working like clock work.
• At MEI, one of the 6 UV GWs is complete, two have the Purcell gap pole tips completely glued down and qualified and one more has its top shims being glued. They plan to ship in a week and one half.
Injector Dipoles (DU/DV)
• No new developments
Arc 180 Degree Dipoles (GY)
• Wang NMR is just about ready to pot the first coil and continues to set up for gluing of Purcell gap shims. They do not want to spoil these operations.
• We will visit them the first two days of next week during these critical operations.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
• Two GQs had their LCW hoses applied by the machinist who is supporting the magnet measurement effort. They are being prepared for measurement at the start of next week. They look good.
• Two GXs will follow.
• Wang continues to vacuum impregnate the new series of coils for the remaining two GX magnets and is converting the mold for the GQ coils to this method.
Trim Quad (QT)
• Trim quads were sent to the FEL for installation and trial fit up. Four will have to be retrofitted with coils to add the GC corrector function to them.
Sextupole (SF)
• Master Machine completed one "Commissioning Sextupole" for trial installation in the enclosure on Monday. The remaining 7 are nearly complete.
Octupole (OT)
• We continue working on the procurement package to get these 2 magnets made in pieces by local shops. We have the coils already.
Beam Line and Vacuum
• Chambers for the arcs continue in production status. We received two chambers this week and two more are due next week. The first arc’s chambers are nearly complete. Master Machine notified us that they would not make their end of November deadline to complete the task for the second arc. The completion will be first week of December. This does not trail the magnet availability schedule, so we should be all right.
• Chicane chamber welding is stopped because of the welder is out sick while chamber machining continues. We may be able to double up on welding production by getting a welder from another group.
• The vacuum pipes between girders leading to the wiggler region were installed.
• The magnet measurement crew stopped installation of the correctors to concentrate on GWs.
• LCW piping fabrication is at 95%.
• The Dump line stand positions (with a variable parameter of the dump face) were issued to the alignment group to start installation. David Douglas is ruminating on the best compromise position for the face that allows the best control of betatron and disbursive beam qualities.
• Layouts of the workarounds above are nearly complete.

WBS 11 (Optics):
The titles of the attached pictures are the best descriptors.
HR vault installation
OC installation

This week's activities culminated with our transporting the optical cavity vacuum vessels to the FEL vault. The support girders with ultraviewers attached, were installed next to their vacuum vessels and aligned by the Survey and Alignment team. A number of incremental milestones, were accomplished to do this, the major ones are listed below:

Working with Survey and Alignment, the internal optical components for both the HR and OC assemblies were fiducialized in the vacuum vessels. We then translated the mirrors for both assemblies, and checked the repeatability. The data reflected the repeatability of the portable CMM, ~ 100 microns. That is our specification. There is evidence that the performance is higher. The HR vacuum vessel, with internal hardware for vertical motion, was leak checked, and checked for oil contamination. It passed both tests. Clean enclosures were established for both the HR and OC assemblies before they were moved.

During the above, we also completed inspection of the first light outcouplers. One failed (visible flaws within 80% CA) but the other passed. We also began installing one of the two alignment HeNe lasers we'll use for our alignment activities.

The custom optical breadboards for the OCMMS optics arrived. Drawings for the insertable mirror used to route high power beam to the diagnostic beam dump are being checked. The vacuum vessels for the high power optical transport are in Procurement.

We discovered that some components for the THz diagnostic still had oil contamination. The components were disassembled, recleaned, and are being baked before they are reassembled.